OpenStudy (anonymous):

5. Solve the equation. Show your work. Check your answer. 4y + 5 = – 31

OpenStudy (godlovesme):

first u have to subtract 5 from both side

OpenStudy (anonymous):

4y + 5 = – 31 4Y=-3-5 4Y=-36/4 =-9 I have this but the teacher said I need to check my work. Iam confused

OpenStudy (anonymous):

subtract ( +5 ) to get 0. So 5-5 is 0, cross it out and the -31 minus 5

Nnesha (nnesha):

@dwatts158 you did a mistake that is not 3 its 31

OpenStudy (anonymous):

that is correct tho

OpenStudy (anonymous):

4Y=-31-5 4Y=-36/4 =-9

Nnesha (nnesha):

there you go :)

OpenStudy (anonymous):

Yeah place -9 into the problem, that is correct :D

OpenStudy (anonymous):

Thats how i check my work

OpenStudy (anonymous):

Good grief... I feel so stupid . lol thanks

Nnesha (nnesha):

nope don't feel that :) O_O

OpenStudy (anonymous):

Yeah, dont feel like that. Most people take longer to understand a question like that xD

OpenStudy (anonymous):

Thank you Nnesha, that is my first medals :D and Dwatts

OpenStudy (anonymous):

And me :P

OpenStudy (anonymous):

Yes xD

Nnesha (nnesha):

you deserve it :) bec you explained very well instead giving direct answer :)